If driving 197 miles . How fast would you have to drive to get there in two hours?
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

d = r*t
r = d/t

Answer by jim_thompson5910(35256)

d = r*t


197 = r*2


197/2 = r


98.5 = r


r = 98.5


You would have to drive 98.5 miles per hour. Definitely not possible (assuming you want to stay within legal limits).